OpenStudy (anonymous):

if anda had 3 hundred dollar bills and 10 50 dollar bills and 6 20 dollar bills will she be able to buy a pair of thebfanciest boots in the country thats $1023 .65 ???

OpenStudy (texaschic101):

3 hundred dollar bills = (3)(100) = 300 10 fifty dollar bills = (10)(50) = 500 6 twenty dollar bills = (6)(20) = 120 now add 300 + 500 + 120 = 920 Now do you think she can spend $1023.65 on those boots ?